OpenStudy (anonymous):

Nancy left a bin outside in her garden to collect rainwater. She notices that 1 over 2 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin please help

OpenStudy (anonymous):

@quickstudent

OpenStudy (anonymous):

if any of you help i will give you a metal and be a fan

OpenStudy (butterflyhope):

1/3 thirds of the water fills 2/3 thirds of the bin

OpenStudy (anonymous):

sorry i didnt post the full question

OpenStudy (anonymous):

@ButterflyHope

OpenStudy (butterflyhope):

Ohhh I was like um that isn't really a question but ok xD

OpenStudy (anonymous):

@jgirl128

OpenStudy (texaschic101):

will a proportion work ? (1/2) / (2/3) = x / (3/3) --- 1/2 gal to 2/3 bin = x gal to 3/3 bin) cross multiply (2/3)(x) = (1/2)(3/3) 2/3x = 3/3 x = (3/3) / (2/3) x = 3/3 * 3/2 x = 9/6 = 1 1/2 the amount of water that would fill the entire bin is : 1 1/2 gallons or 1.5 gallons